OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. Question 5 options: 1) 203 2) 197 3) 168 4) 163

OpenStudy (angela210793):

a18=a1+17*d 95=-7+17d 17d=102 d=6 a35=a1+(n-1)*d=-7+34*6=197